OpenStudy (anonymous):

if the fifth term of an arithmetic sequence is -5 and the ninth term is -17, the first term of the sequence must be a. -7 b. -5 c. -11 d. 7

OpenStudy (campbell_st):

well I suppose you need to set up the equations 5th term \[-5 = a_{1} + (5 -1)d~~~~or~~~-5 = a_{1} + 4d\] 9th term \[-17 = a_{1} + (9 -1)d~~~~~or~~~~~-17 = a_{1} + 8d\] so the 2 equations are -5 = a + 4d -17 = a + 8d subtract the equations to find d then substitute the value into either equation to find a